Sol Summary – March 17th

Crew 309 Sol Summary Report 17Mar2025

Sol: 1

Summary Title: First day of work

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • This morning Co-Commander A and Scientist completed an EVA while the remainder of the crew worked on tasks in the hab. In the afternoon, Engineer and Medic completed an EVA, followed by a final EVA completed by Co-Commander B and GreenHab.
  • The team is adjusting to life on Mars. We encountered a number of technical issues and are working to learn our crew responsibilities. Despite these challenges, the team remains optimistic.

Look Ahead Plan: We will be conducting two EVAs tomorrow in the morning and afternoon

Anomalies in work:

  • Solution for eye tracking device 681 is to use an external battery pack while they are on EVA so that it doesn’t deplete and turn off. The MC external battery packs need a lot of time to charge, so we are using Jesus.
  • This is not a possibility for the device 691 because it does not have an external charging port. The solution for this one is to add the battery as late as possible. The battery packs are dead right now, but Jesus is using his personal.
  • The depressurization door opened, need to close it more securely.

Weather: Sunny and windy, 43F / 70F

Crew Physical Status: Nominal

EVA: 3 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report,

Support Requested: None

Sol Summary – March 18th

Crew 309 Sol Summary Report 18Mar2025

Sol: 2

Summary Title:

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • This morning commander A and crew engineer went on their 2nd EVA of the week. Today was especially windy and so the crew was mainly upstairs to keep warm because the heaters were struggling to keep up. Following lunch the crew noticed the smell of propane down stairs and began investigating the heater. We determined that the high winds were affecting the heating system. The high winds also knocked down a safety ladder which the crew had to secure. Eventually we had to shut off the heater for a few hours and just use a small space heater upstairs. As the winds began to slow we were able to turn the heat back on and our second EVA, which was delayed by the weather, was completed successfully. In the evening we had a good dinner and conversed about lessons learned from the day.

Look Ahead Plan: We will be conducting two EVAs tomorrow in the morning and afternoon

Anomalies in work:

  • The heating system was set to 70 and then 72 but the hab stayed consistently at 67. We believe this is due to the strong winds. An hour later it has risen to 68 but has not reached 72. Upon examination, the heating system is not consistently blowing hot air. Current time is about 10:00 AM.
  • Crew smelled a propane smell, and a gas alarm went off. Engineer used a gas detector and no leaks were detected. Crew examined the HVAC system and found an error code blinking. The diagnosis is that the excessive winds are affecting the pressure switch of the HVAC causing it to turn on and off. This diagnosis explains the HVAC not staying continuously on until target temperature is reached and the propane smell. Current temperature is set to reach 72 degrees, but the temperature is at 67. Winds are at about 30 mph and should be less tonight. Current time is about 1:00 PM.
  • Crew communicated with Mission Support to replace the HVAC filter. Crew has replaced the air filter for the HVAC. Current time is about 2:45 PM.
  • The propane smell persisted and became substantial at 3:00 PM. We are turning the HVAC off and using the space heater on the green/low setting.
  • Around 5:30 PM, the wind reduced and the crew turned the HVAC system back on. We have experienced and has had no issues going

Weather: Windy, 27F / 48F

Crew Physical Status: Nominal

EVA: 2 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report

Support Requested: None

Sol Summary – March 19th

Crew 309 Sol Summary Report 19Mar2025

Sol: 3

Summary Title:

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • Commander B and crew engineer started the day by performing and EVA. The crew medic continued work on our robot meant to transport medical supplies. The crew is continuing to experience issues with the sensors we’re supposed to wear during EVAs so the crew scientist and engineer began troubleshooting. In the afternoon the GreenHab officer and Medic completed an EVA.

Look Ahead Plan: We will be conducting two EVAs tomorrow in the morning and afternoon

Anomalies in work:

  • The Tobbii eye-tracking glasses disconnected after 30 minutes on EVA 6. Engineer and Scientist are troubleshooting why this may be. Only theory is that the glasses occasionally stop recording when they are out of range of the computer’s Wi-Fi.
  • Fuzzy part of the mic was having difficulty staying attached to the microphone. It fell off mid-EVA but EVA 6 was able to find the fuzzy component and bring it back to the Hab.
  • For EVA 7, we are experimenting with hooking up the eye-tracking glasses with the external battery to the portable laptop to see if we can extend eye tracking recording for at least one participant.
  • Update: the heating issue is resolved now that there are no longer excessive winds.
  • Accidentally made 6x the amount of onions than intended. Read instructions in the future.

Weather: Sunny, 53F / 29F

Crew Physical Status: Nominal

EVA: 2 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report

Support Requested: None

Sol Summary – March 20th

Crew 309 Sol Summary Report 20Mar2025

Sol: 4

Summary Title:

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • Commander A and B went on an EVA in the morning. The crew scientist and engineer continued troubleshooting the eye tracking equipment and think they have found a solution, but they are unable to implement it due to a lack of an ethernet cable on Mars. The entire crew continued working on the mission vlog throughout the day. In the afternoon the Greenhab and science officer went on and EVA. Martian activity was high today. A malfunction of the wifi system delayed crew reports by 2+ hours.

Look Ahead Plan: We will be conducting three EVAs tomorrow in the morning and afternoon

Anomalies in work:

  • The Tobbii eye tracking glasses are now not connecting to either computer at all. Scientist is troubleshooting.
  • We found a potential default setting which may be shutting the Tobii controllers down when it starts to go low on the battery. An ethernet cable needs to be used to connect the controllers. The app may also be out of date. We also were able to reconnect the glasses, the default password needed to be used.
  • We tried to use the Tobii glasses without an internal battery but just with the external battery, this approach was not effective.

Weather: Windy, 61F / 26F

Crew Physical Status: Nominal

EVA: 2 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report

Support Requested: None

Sol Summary – March 21st

Crew 309 Sol Summary Report 21Mar2025

Sol: 5

Summary Title: Final Day

Author’s name: Co-Commander A

Mission Status: Nominal

Sol Activity Summary:

  • We conducted 3 EVAs today. GreenHab and Engineer went in the morning, commander A and Medic in the afternoon, and commander B and Scientist last. The crew began the cleaning process completing the RAM, GreenHab, and Food Inventory tasks this evening. Simulation ended after dinner and the crew was able to explore some of the Martian terrain outside of our EVA suits.

Look Ahead Plan: We will be packing up and heading home in the morning.

Anomalies in work:

  • During prep for our second EVA of the day, we determined that the helmet had become detached from the spacesuit worn by crew member one. Engineer, Medic, and Commander A were able to repair the suit for the remainder of the EVAs. A more permanent solution will be required for future missions.

Weather: Windy, 34F / 62F

Crew Physical Status: Nominal

EVA: 3 EVAs completed

Reports to be filed: Sol Summary, Journalist Report, GreenHab, Operations Report, Mission Summary

Support Requested: None

Sol Summary – February 18th

Crew 311 Sol Summary Report 18-02-2025
Sol: 2
Summary Title: “Why do astronauts have laundry problems? Because they are always in space!”
Author’s name: Erin Pougheon, Crew Commander
Mission Status: Nominal
Sol Activity Summary: SOL 2 started nominally for crew 311 with Core Data retrieval and sport session, organized by our Crew HSO Célyan.
After that, Crew Journalist, Crew Engineer and Crew Astronomer finished building the atmospheric instruments in the RAM for the next EVA, during which they’ll install them on Mars. Meddi, our GreenHab Officier, continued working in the GreenHab. Some crew members also conducted some cognitive assessments for the Orbital Architecture experiment.
During the afternoon, Crew Astronomer did dome solar observations in the Observatory, taking advantage of the nice weather. Meanwhile, Crew Commander worked on the crew schedule for the next SOLs and other crew members worked on installing other experiments, such as AMAIA, the AI assistant for astronauts which still has some humor problems! Crew Engineer and Crew HSO started installing the light in the Hab for the Lattal experiment.
The evening is calm, and every crew member is filling their questionnaires before the beginning of the communication window.
Look Ahead Plan: Tomorrow will be centered on the EVA of installation for the atmospheric instruments. For crew members not on EVA, or during the afternoon, installing the experiment and testing that everything is working nominally will be a priority.
Anomalies in work: None
Weather: Sunny, Temperature 40 F / 50 F
Crew Physical Status: Nominal
EVA: None
Reports to be filed: Sol Summary, EVA Request #03, Operations Report, GreenHab Report, Astronomy Report, Journalist Report, Crew Photos.
Support Requested: None

Sol Summary – February 19th

Crew 311 Sol Summary Report 19-02-2025
Sol: 3
Summary Title: A break to freshly restart the day!
Author’s name: Erin Pougheon, Crew Commander
Mission Status: Nominal
Sol Activity Summary: This morning, three crew members went on EVA to install 4 atmospheric instruments: LOAC, which analyze the atmosphere and aerosols, MegaAres and the Field Mill, which analyze the electric field and a weather station. The EVA went very well, and they even finished earlier than planned. While the EVA was going on, Crew Scientist started the HUMANISE experiment by teleoperating a rover.
After a delicious lunch prepared by Crew HSO and Crew Scientist, the whole crew took a well-deserved break. The afternoon was quite calm but productive: GreenHab Officer finished planting the micropouss and took photos for our outreach activities!
As every day, the end of the afternoon is fully dedicated to filling in questionnaires and preparing the communication window!
Look Ahead Plan: Tomorrow will be dedicated to continuing the installation of the experiments we started, such as testing that all the data from the Orbital Architecture experiment is well collected. We also want to start working on new experiments, such as for example the building of a rover from scratch with a 3D printer.
Anomalies in work: None
Weather: Cloudy, Temperature 30 F / 50 F
Crew Physical Status: Nominal
EVA: None
Reports to be filed: Sol Summary, EVA Report #03, Operations Report, GreenHab Report, Journalist Report, Crew Photos.
Support Requested: None

Sol Summary – February 20th

Crew 311 Sol Summary Report 20-02-2025
Sol: 4
Summary Title: The wind blows in the sails of Crew 311
Author’s name: Erin Pougheon, Crew Commander
Mission Status: Nominal
Sol Activity Summary: Sol 4 allowed Crew 311 to continue installing some experiments, and to start some sessions of the HUMANISE experiment. Crew Engineer started teaching 3D printing to other crew members. We tried to debug AMAIA, our AI assistant, and it works but not perfectly. We also launched for the first time all the measurements for the Orbital Architecture experiment (heart rate sensor, location tracking,…).
Crew Astronomer and GreenHab Officer thought about how to organize the EVA that will take place during Sol 6. The objective will be to take some drone pictures for the stratigraphy experiment.
During the afternoon, Crew continued their activities and finished installing the lights in the Lower Deck for the Lattal experiment.
Look Ahead Plan: Tomorrow will be dedicated to the installation of some experiments. Some crew members will create a sign to place near the atmospheric instruments during the next EVA and other features to monitor the battery. We’ll also start to count the water used and for which usage. GreenHab Officer will also train some other crew members on how to use the drone.
Anomalies in work: None
Weather: Cloudy, Temperature 35 F / 45 F
Crew Physical Status: Nominal
EVA: None
Reports to be filed: Sol Summary, Operations Report, GreenHab Report, Astronomy Report, Journalist Report, Crew Photos.
Support Requested: None

Sol Summary – March 12th

Crew 311 Sol Summary Report 12-03-2025

Sol: 24

Summary Title:

Author’s name: Erin Pougheon, Crew Commander

Mission Status: Nominal

Sol Activity Summary: This morning, we performed the last EVA of photogrammetry at Kissing Camel Ridge W. GreenHab Officer and Crew Journalist successfully found all 10 checkpoints in the area! Meanwhile, in the Hab, activities and experiments continued to happen as planned.

For lunch, we ate delicious Marian-made burgers, it was incredible!

In the afternoon, we had our last session of brainstorming on how to improve the station, for the Orbital Architecture experiment. It was about the Science Dome, the Observatory and the whole station organization, such as the tunnels and module disposition.

Look Ahead Plan: Tomorrow, we’ll stay at the station, continuing the experiments we have done the whole mission. In the afternoon, we’ll start to uninstall and pack some of them, such as Lattal or the 3D printer.

Anomalies in work: None

Weather: Sunny, Temperature 45 F / 60 F

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, EVA Report #15, Operations Report, GreenHab Report, Journalist Report, Crew Photos.

Support Requested: Do we need to send a Research/Scientific report at the end of the mission (sol 26) as we did at mid-mission?

Sol Summary – March 13th

Crew 311 Sol Summary Report 13-03-2025

Sol: 25

Summary Title:

Author’s name: Erin Pougheon, Crew Commander

Mission Status: Nominal

Sol Activity Summary: For our last full day of mission, we worked a lot! With no EVA in the morning, we had plenty of time to perform the last sessions of experiments, such as the VR for HUMANISE. We also had time to finish documentation for our Core Data and Crew Astronomer finished his code for the astronomy project!

In the afternoon, we filmed some videos for outreach purposes and continued the VR sessions. We also started to uninstall and pack some experiments, such as the Lattal’s lights and the 3D printer.

Look Ahead Plan: Tomorrow, we’ll perform an EVA in the morning to pick up the atmospheric instruments that are at Marble Rituals. Then, we’ll pack them, such as every other experiment material we have with us at the station. Meanwhile, the last sessions for Orbital Architecture will continue to run in parallel. We’ll break the simulation at 5:00 PM, by leaving the station!

Anomalies in work: None

Weather: Cloudy, Temperature 50 F / 55 F

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, EVA Request #16, Operations Report, GreenHab Report, Journalist Report, Crew Photos.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site